### v3

# Collision Probability Estimation for Optimization-based Vehicular Motion Planning

Abstract
Many motion planning algorithms for automated driving require estimating the probability of collision (POC) to account for uncertainties in the measurement and estimation of the motion of road users. Common POC estimation techniques often utilize sampling-based methods that suffer from computational inefficiency and a non-deterministic estimation, i.e., each estimation result for the same inputs is slightly different. In contrast, optimization-based motion planning algorithms require computationally efficient POC estimation, ideally using deterministic estimation, such that typical optimization algorithms for motion planning retain feasibility. Estimating the POC analytically, however, is challenging because it depends on understanding the collision conditions (e.g., vehicle’s shape) and characterizing the uncertainty in motion prediction. In this paper, we propose an approach in which we estimate the POC between two vehicles by over-approximating their shapes by a multi-circular shape approximation. The position and heading of the predicted vehicle are modelled as random variables, contrasting with the literature, where the heading angle is often neglected. We guarantee that the provided POC is an over-approximation, which is essential in providing safety guarantees. For the particular case of Gaussian uncertainty in the position and heading, we present a computationally efficient algorithm for computing the POC estimate. This algorithm is then used in a path-following stochastic model predictive controller (SMPC) for motion planning. With the proposed algorithm, the SMPC generates reproducible trajectories while the controller retains its feasibility in the presented test cases and demonstrates the ability to handle varying levels of uncertainty.
Index Terms:
Collision Probability, Circular Approximations, Stochastic Model Predictive Control, Path-following
I Introduction
Automated vehicles (AVs) usually operate with limited knowledge about their surroundings, which stems from limitations in sensing and estimation, e.g., range limitations or physical sight obstructions, noise in the sensed data, or missing information about the intentions of other road users. This limited knowledge, commonly referred to as uncertainty, is an important factor that should be incorporated in the design of an AV’s motion planning and control algorithms as recommended in scientific literature (see, e.g., [schwarting2018planning, McAllister.82017, hubmann2018automated, Tolksdorf_2025]) and safety standards [ISO21448].
The specifics of how to incorporate uncertainty in the design, however, depends on the particular algorithms and their implementation. Regarding motion planning and decision-making, uncertainty in the position and orientation of other actors and objects is understood as a challenge to ensure safety. Here, the probability of collision (POC) [schwarting2017safe, goulet2022probabilistic, schreier2016integrated, althoff2009model] or related quantities, e.g., stochastic risk [tolksdorf2023risk, hruschka2019uncertainty, nyberg2021risk], are measures commonly used to ensure safety by constraining or minimizing them within a motion planning algorithm.
As mentioned, from an AV perspective, the positions and heading angles of other vehicles are uncertain. A common method to express this uncertainty is by treating the uncertain variables as random variables with specific associated probability density functions (PDFs). The POC for two arbitrarily shaped vehicles is then typically estimated with Monte Carlo sampling (MCS) [lambert2008collision, schreier2016integrated]. That is, the POC is estimated by sampling the positions and orientations from their PDFs and checking whether these positions and orientations lead to a collision. The ratio of colliding to non-colliding samples then gives the POC estimate. We will refer to this method for POC estimation as MCS. Although applicable for any vehicle shape, MCS is computationally expensive as many samples are necessary for an accurate estimation; additionally, the POC estimate fluctuates around its true value for each sampling set. Hence, an inherent MCS drawback is that it is not guaranteed to not under-approximate the POC for a finite amount of samples, which is generally undesirable in safety-critical applications.
To improve upon the MCS approach, various analytic methods for POC estimation have been proposed. For instance, using occupancy grids. That is, the motion planning space in $\mathbb{R}^{2}$ is discretized into cells, and the probability of occupation by both actors of a cell is calculated [althoff2011comparison], [candela2023risk]. In [althoff2011comparison], the POC is approximated with Markov chains and compared with MCS, where it is concluded that MCS is the better approximation in accuracy and computational efficiency. Aside from occupancy grids, [du2011probabilistic] chooses to model a robot and an obstacle as a circle each and formulates a joint Gaussian PDF of each vehicle’s position to approximate the POC analytically. The authors of [philipp2019analytic] consider rectangular-shaped vehicles, of which the position of one is known, whereas the position of the other vehicle is Gaussian distributed. From the distributions, the PDFs are numerically integrated to retrieve the POC; however, both vehicles’ heading angles are assumed to be deterministic. The same approach, i.e., using rectangles with deterministic heading angles, is utilized by [altendorfer2021new] to formulate the POC over a time interval, where an analytic solution is derived for an approximation of the POC. An analytic approximation of the POC is also proposed by [patil2012estimating]; here, the vehicle’s shape is assumed to be point-like and the uncertainty is Gaussian.
Another shape approximation in motion planning is the multi-circular approximation, where a rectangle is covered with overlapping circles [ziegler2010fast, ziegler2014trajectory, schwarting2017safe, gutjahr2016lateral, manzinger2020using, werling2012optimal, micheli2023nmpc]. Computing the POC analytically for a multi-circular AV approximation has been presented in [tolksdorf2024], where it is also shown that the proposed analytic method significantly outperforms MCS in computational speed while a slight over-approximation of the true POC is guaranteed. In spite of the benefits, the proposed method only considers the AV to be approximated by multi-circles and all vehicles with whom it is potentially colliding by a single circle, avoiding the need to consider the heading angle, yet undermining the accuracy of this approximation. In [mustafa2024racp], the vehicle’s shapes are each approximated by a two-circle covering, though the heading angle is assumed to be deterministic. Evidently, for estimating the POC for two vehicles of which the position and orientation of one vehicle is uncertain, the main challenge is to guarantee not to underestimate the true POC while generating a not-too-conservative overestimation.
The motion plan of an AV can be generated with any of several techniques (see, e.g., [Gonzalez.2016, Claussmann.2020]). Here, we focus on stochastic model predictive control (SMPC) as an optimization-based technique. SMPC allows for the flexible integration of the POC either as a constraint or within a cost function to be minimized. Further, path-following SMPC (see [Faulwasser.2016] for details) has been shown to generate smooth trajectories while successfully adapting the motion plan to a given level of uncertainty [tolksdorf2023risk]. The drawbacks of SMPC, however, are its high computational burden and the general lack of guarantee that it is recursively feasible. Further, the evaluation of probabilistic constraints (e.g., constraints based on the POC) is known to be challenging, as, in many cases, sampling-based estimation techniques (online MCS) are computationally heavy or analytic methods are strongly over-approximative, leading to overly conservative behavior [mesbah2016stochastic]. Suppose, for instance, the POC is estimated with MCS within an optimization-based motion planner. If the optimizer converges towards the boundary of the POC constraint, a previously feasible point might suddenly become infeasible due to an unfavorable random draw [homem2014monte]. Therefore, results, i.e., planned trajectories, are non-reproducible as the fluctuations in POC estimation are random, and the optimizer might struggle to converge as a consequence.
The main contributions of this paper are as follows: we address the aforementioned challenges for computing the POC by presenting a novel method based on multi-circular shape approximations for all actors to estimate the POC, particularly for motion-planning applications. This method overcomes the drawbacks of MCS, i.e., its high computational load, the lack of a guarantee to over-approximate the POC, and POC fluctuations causing the planning optimizers to fail. We present the necessary geometric conditions for multi-circle collisions to guarantee that the method provides an over-approximation, while practical case studies demonstrate that the results are not overly conservative. Furthermore, we present an algorithm to compute the POC for a Gaussian positional uncertainty and a wrapped Gaussian distribution for the heading angle. On the basis of a representative numerical case study, we discuss the trade-off between computational efficiency and over-approximation accuracy. Lastly, we present an SMPC scheme for motion planning leveraging the advantages of the proposed POC estimation method: computing smooth, reproducible trajectories while the controller retains its recursive feasibility in the tested scenarios and reacts conservatively when the uncertainty grows.
The remainder of this article is organized as follows. Section II provides preliminary notation and presents the problem statement of computing the POC using multi-circle shape approximations. Further, Section III bounds the set of all collision configurations. Section LABEL:sec_combination_of_proabilities guarantees the over-approximation of the POC. Section LABEL:sec_gaussian_POC_algo presents a method to compute the POC for Gaussian uncertainty, and the SMPC motion planning scheme is presented in Section LABEL:sec_SMPC. The POC estimation for Gaussian uncertainties and the SMPC are then simulated in representative case studies in Section LABEL:sec_simulation_example, and the results are discussed in Section LABEL:sec_discission. Finally, conclusions and recommendations for future work are given in Section LABEL:sec_conclusions_and_future_work.
II Problem Statement: POC Estimation
Figure 1: Schematic of the problem statement.
Consider an arbitrary traffic scene consisting of an automated vehicle and another vehicle, referred to as ego and object vehicles. As the ego performs the POC estimation, all variables are defined with respect to the ego’s geometric center. By assumption, the object does not communicate any information to the ego (e.g., its position or intentions). We characterize each vehicle by a configuration composed of a position and orientation. Let any position in $\mathbb{R}^{2}$ be denoted by $\mathbf{x}:=(x,y)$ and any orientation be denoted by $\theta\in\mathbb{R}_{\geq 0}$. Together, they construct the configuration space $\mathcal{C}:=\mathbb{R}^{2}\times\mathbb{R}_{\geq 0}$, giving the set of all possible vehicle configurations. For later calculations, we will restrict $\theta$ to positive values for convenience. Whenever confusion can arise, we identify variables associated with the ego and the object, respectively, with $e$ and $o$ subscripts. Let $\mathbf{y}_{o}:=(\mathbf{x}_{o},\theta_{o})\in\mathcal{C}$ characterize the object’s configuration. That is, $\mathbf{x}_{o}$ denotes the position of its geometrical center and $\theta_{o}$ denotes its heading angle measured with respect to the ego’s $x$-axis (see Figure 1). A set of integers $\{a,a+1,...,b\}$, with $a<b$, is denoted by $\mathbb{Z}_{a}^{b}$. We denote a configuration $\mathbf{y}$ at discrete-time instant $k$ as $\mathbf{y}_{k}$ and a predicted configuration at discrete-time instant $n\geq k$ given information available at time $k$ as $\mathbf{y}_{n|k}$. The same notation is applied for states, inputs, and constraints. All random variables are defined over the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ is a sigma-algebra over $\Omega$, and $\mathbb{P}$ is a probability measure over $\mathcal{F}$. The sets in $\mathcal{F}$, also known as events, are denoted by $\mathcal{A}_{m}$ with $m\in\mathbb{N}$. In the sequel, the following definitions will be needed.
Definition 1.
(Probability Measure) A mapping $\mathbb{P}:\mathcal{F}\rightarrow[0,1]$ is a probability measure if:
  • (i)
    $0\leq\mathbb{P}\{\mathcal{A}_{m}\}\leq 1$, for any $\mathcal{A}_{m}\in\mathcal{F}$,
  • (ii)
    $\mathbb{P}\{\Omega\}=1$ and $\mathbb{P}\{\emptyset\}=0$,
  • (iii)
    $\mathbb{P}\{\bigcup_{m=1}^{\infty}\mathcal{A}_{m}\}=\sum_{m=1}^{\infty}\mathbb{P}\{\mathcal{A}_{m}\}$, if $\mathcal{A}_{m}$ are pairwise disjoint (i.e., $\mathcal{A}_{m}\cap\mathcal{A}_{j}=\emptyset$, for $m\!\neq\!j$ and $\mathcal{A}_{m},\mathcal{A}_{j}\in\mathcal{F}$).
Definition 2.
[…]
II-A Collision Probability for Arbitrary Shapes
Figure 2: The ego estimates and predicts the uncertainty about the object’s current and future configurations. Note, the PDFs $p_{\mathbf{y}_{o,n|k}}$ may have a different functional form for each time-step, thus accommodating different possible object behaviors.
For motion planning purposes, the ego vehicle (referred to as ego from now on) must estimate the POC with the object vehicle (referred to as object from now on) at all times along its motion plan. Typically, the ego is equipped with a sensor suit that perceives the object, from which the information is processed by algorithms estimating its current configuration and predicting the future configurations of the object. Clearly, the last step, i.e., predicting future configurations of the object, introduces the greatest amount of uncertainty, especially given an already uncertain current configuration estimate. Hence, we abstract the process of measurement, estimation, and prediction by letting the object’s configuration be uncertain at any point in time. That is, at a current time instance $k$, or predicted time instance $n\geq k$, a PDF can characterize the uncertainty in the configuration of the object. We note that this concept can be applied to many state-of-the-art road user prediction models, where the PDF can be any distribution over the configuration space $\mathcal{C}$, even accommodating very different object behaviors (see the example illustrated in Figure 2). Consequently, we adopt the following assumption about the object.
Assumption 1.
(Object information) The ego estimates the kinematic variables $\mathbf{y}_{o,n|k}$ at discrete-time instance $n\geq k$, given information available at time $k$, together with their associated probability density function $p_{\mathbf{y}_{o},n|k}$, for all instances $n$ within a prediction horizon $\mathbb{Z}_{k}^{k+N_{P}}$. All components in $p_{\mathbf{y}_{o},n|k}$ are mutually independent.
A collision occurs if the spaces occupied (i.e., the footprints) by the ego and the object intersect, regardless of how both got there. Thus, we note that the velocity is not considered to determine the POC, as the configuration is evidently a consequence of applying a velocity over time [^1] Consider, e.g., that the object configurations are predicted by a dynamic model, propagating estimated object inputs (steering and acceleration) to object configurations. Here, the inputs are random variables, causing the predicted configurations also to be random variables. However, only the configurations are required to determine the occurrence of collisions, and, hence, the POC estimation method is independent of the object’s inputs. Nonetheless, the object’s inputs are implicitly considered by the prediction model, propagating those into (future) configurations; thus, the estimation result is dependent on the object’s inputs.. Given the fact that the method of POC estimation is the same for all time instants, we omit time indexing in the sequel for notational simplicity. Without loss of generality, we assume that both footprints are rectangles over $\mathbb{R}^{2}$ denoted, respectively, by $\mathcal{S}_{e}$ and $\mathcal{S}_{o}(\mathbf{y}_{o})$ (see dashed boxes in Figure 1). As all measurements are with respect to the ego, the POC is a function of the statistics of $\mathbf{y}_{o}$ only. More specifically, if $\tilde{\mathcal{A}}_{rec}:=\{\mathbf{y}\in\mathcal{C}\mid\mathcal{S}_{e}\cap\mathcal{S}_{o}(\mathbf{y})\neq\emptyset\}$ denotes the object configurations that lead to a collision, then the POC is given by
$\textup{POC}\triangleq\mathbb{P}\{\mathbf{y}_{o}\in\tilde{\mathcal{A}}_{rec}\}=\int\displaylimits_{\tilde{\mathcal{A}}_{rec}}p_{\mathbf{y}_{o}}(\mathbf{y})\text{d}\mathbf{y}.$ (1)
II-B Multi-Circle Footprint Approximations
Although both footprints $\mathcal{S}_{e}$ and $\mathcal{S}_{o}(\mathbf{y}_{o})$ are considered rectangular, we propose to use a multi-circle footprint cover. In the sequel, $\mathcal{B}[\mathbf{x}_{c};r]$ denotes a closed circle with center $\mathbf{x}_{c}=(x_{c},y_{c})$ and radius $r\in\mathbb{R}_{>0}$. That is,
$\mathcal{B}[\mathbf{x}_{c};r]:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\|\mathbf{x}_{c}-\mathbf{x}\|\leq r\}.$ (2)
Suppose that the smallest rectangular approximation of a vehicle’s footprint has length $l$ and width $w$, $l\geq w$. Further, suppose a vehicle’s footprint is covered by $N_{c}$ overlapping, closed circles with equal radii $r$ and centers $\mathbf{x}_{i}(\mathbf{y})$, $i\in\mathbb{Z}_{1}^{N_{c}}$, where $\mathbf{y}$ denotes the vehicle configuration. That is,
$\mathcal{S}(\mathbf{y})\subset\mathcal{BC}(\mathbf{y}):=\bigcup_{i=1}^{N_{c}}\mathcal{B}[\mathbf{x}_{i}(\mathbf{y});r].$ (3)
It can be shown that the smallest radius $r$ needed to fully cover the rectangle, when all circles are placed equidistantly along the central longitudinal axis of the rectangle (with distance $d_{c}$), is given by
$r\!=\!\sqrt{\left(\!\frac{l}{2N_{c}}\!\right)^{2}\!+\!\frac{w^{2}}{4}},\;d_{c}\!:=\!2\sqrt{r^{2}\!-\!\frac{w^{2}}{4}}\!=\!\|\mathbf{x}_{i+1}(\mathbf{y})\!-\!\mathbf{x}_{i}(\mathbf{y})\|$ (4)
with $i\in\mathbb{Z}_{1}^{N_{c}-1}$. The problem setup is depicted in Figure 1 for a two-circle ($N_{c}=2$) vehicle footprint covering.
Problem.
Derive an estimate of the POC in (1) by using the ego and object footprint covering (2) - (4).
The problem is approached by first identifying the intersection conditions (i.e., conditions when multi-circular coverings of two vehicles collide) for multi-circular coverings, yielding a set of all possible intersection configurations [^2] As will be shown in Theorem LABEL:theo_over_approx, the collision configurations are a subset of the intersection configurations presented in Section III, which is why our POC estimate is guaranteed to be an over-approximation of the actual POC., and subsequently, integrating the PDF $p_{\mathbf{y}_{o}}$ over this set to obtain the POC estimate.
III Intersection Conditions for Multi-Circle-to-Multi-Circle Collisions
Figure 3: (a) Problem setup for some $(\phi_{o},\rho_{o},\theta_{o})$ and shifting the polar frame by distance $L_{e}$ to $\rho^{\prime},\phi^{\prime}$. (b) Upper-bounding $\rho^{\prime}$. (c) Bounding $\theta_{o}$ for a given $\rho^{\prime}$ and $\phi^{\prime}$.
To estimate the POC from the given PDF $p_{\mathbf{y}_{o}}$, one is tasked with defining a set $\tilde{\mathcal{A}}\subset\mathcal{C}$, for which conditions for a collision are satisfied, see (1). Clearly, some object configurations can be excluded as they are too distant from the ego for a collision to occur, irrespective of the object’s heading angle. In fact, let us define the relative distance of the object’s geometric center to the ego’s geometric center as $\rho_{o}=\sqrt{x_{o}^{2}+y_{o}^{2}}$. Further, denote $R:=r_{e}+r_{o}$ as the joint radius, with $r_{e}$ and $r_{o}$, respectively, the radii of the circles covering the ego and the object (given by (4)). Herewith, we can upper-bound the relative distance for which a collision can occur to a maximum collision distance as
$\overline{\rho}=R+\frac{d_{c,o}}{2}\big(N_{c,o}-1\big)+\frac{d_{c,e}}{2}\big(N_{c,e}-1\big),$ (5)
where $N_{c,e},N_{c,o}$ denote the number of circles covering the ego and object and $d_{c,e},d_{c,o}$ are defined as in (4) for the ego and object, respectively. Note that for any relative distance $\rho_{o}>\overline{\rho}$, a collision cannot occur. Using this inequality and the radial bounding (5) intuitively motivates the usage of a polar coordinate frame, where we define a change of coordinates as
$\begin{split}&CT:\mathbb{R}\times[0,2\pi)\rightarrow\mathbb{R}^{2},\\ &(\rho,\phi)\mapsto(x,y)=(\rho\cos{\phi},\rho\sin{\phi}).\end{split}$ (6)
Using (6) and the two-argument arctangent function ‘atan2’, the random vector
$\mathbf{y}_{o,p}:=(\phi_{o},\rho_{o},\theta_{o})=(\text{atan2}(y_{o},x_{o}),\sqrt{x_{o}^{2}+y_{o}^{2}},\theta_{o})$ (7)
is derived, where the subscript $p$ denotes that $\mathbf{y}_{o}$ is expressed in polar coordinates [^3] Note that $\mathbf{y}_{o,p}=(\mathbf{x}_{o,p},\theta_{o})$; where $\mathbf{x}_{o,p}=CT^{-1}(\mathbf{x}_{o})$, hence, $\mathbf{x}_{o,p}=(\phi_{o},\theta_{o})$ is no longer Gaussian, and $\phi_{o}$ and $\rho_{o}$ are in general not independent (although $\mathbf{x}_{o,p}$ and $\theta_{o}$ remain independent).. As such, we can restrict the position coordinates of object positions leading to a collision to $(\phi_{o},\rho_{o})\in[0,2\pi)\times[0,\overline{\rho}]$, since outside this set, no collisions can occur, irrespective of the object’s heading. Within that set, however, several conditions for collisions can be found. To examine those conditions, we first study the base case of two offset-circles intersecting. The more general collision conditions are derived later based on this base case.
III-A Base Case: Intersection Conditions for Offset-Circles
We denote circles as offset when their centers are not placed in the geometric center of their associated vehicles but offset by a distance $L_{e}$ and $L_{o}$. The base case is depicted in Figure 3(a). As the ego circle is offset by the distance $L_{e}$ along the ego’s $x$-axis, we transform $(\phi_{o},\rho_{o})$ to the offset center. The transformed polar coordinates are obtained as
$\phi^{\prime}=\text{atan2}(y_{o},(x_{o}-L_{e}))\text{ and }\rho^{\prime}=\sqrt{(x_{o}-L_{e})^{2}+y_{o}^{2}}.$ (8)
Note that $x_{o},y_{o}$ are related to $\phi_{o}$ and $\rho_{o}$ by (6). Evidently, for an intersection to occur in the shifted polar coordinate frame, a new radial upper-bound $\overline{\rho}^{\prime}:=R+L_{o}$ on $\rho^{\prime}$ is found, see Figure 3(b). Whenever the intersection depends on the heading angle $\theta_{o}$, the lower-bound $\underline{\theta}$ and upper-bound $\overline{\theta}$ characterize the first and last contact point between both circles (see Figure 3 (c)), which can be determined geometrically as
$\begin{split}&\overline{\theta}(\phi^{\prime},\rho^{\prime})=\phi^{\prime}+\pi+\arccos{\left(\frac{L_{o}^{2}+\rho^{\prime 2}-R^{2}}{2L_{o}\rho^{\prime}}\right)},\\ &\underline{\theta}(\phi^{\prime},\rho^{\prime})=\phi^{\prime}+\pi-\arccos{\left(\frac{L_{o}^{2}+\rho^{\prime 2}-R^{2}}{2L_{o}\rho^{\prime}}\right)}.\end{split}$ (9)
The full derivation of (9) is provided in Appendix LABEL:appendixA. The conditions for offset-circle intersections are summarized in Table III-A. Here, the column Intersection angle interval gives the range of heading angles that lead to an intersection under the condition that the Radial condition is satisfied. An empty interval (i.e., $\emptyset$) indicates that the circles cannot intersect for any angle $\theta_{o}\in[0,2\pi)$.
TABLE I: Intersection Conditions, Base case.
\__nicematrix_patch_booktabs:
\__nicematrix_revert_colortbl: